# Sparsity-Promoting Reachability Analysis and Optimization of Constrained Zonotopes

### v1

@@ VI-C Moving Horizon Estimation and Set-Valued State Estimation
Consider the combined MHE and SVSE problem described in Sec. V-B with a noise-perturbed double integrator model where $A$, $B$, $\mathbf{x}_{k}$, and $\mathbf{u}_{k}$ are defined as in (34) and have the same dimensions as in Sec. VI-B. The discrete time step is $\Delta t=1$ s. The set of possible states is given as $\mathcal{S}=\mathcal{S}_{p}\times\mathcal{S}_{v}$ where $\mathcal{S}_{p}$ and $\mathcal{S}_{v}$ are both regular hexagons. The inner radii of $\mathcal{S}_{p}$ and $\mathcal{S}_{v}$ are 500 m and 1 m/s, respectively. The process and measurement noise similarly have the forms $\mathcal{W}=\mathcal{W}_{p}\times\mathcal{W}_{v}$ and $\mathcal{V}=\mathcal{V}_{p}\times\mathcal{V}_{v}$ where $\mathcal{W}_{p}$, $\mathcal{V}_{p}$, $\mathcal{W}_{v}$, and $\mathcal{V}_{v}$ are regular hexagon approximations of circular bounds on position and velocity noise distributions. Truncated normal distributions are used such that the position and velocity noise sets have inner radii of $2\sigma$. Zero-mean noise is used in all cases. The position process noise has standard deviation $\sigma=0.001$ m, and the velocity process noise has standard deviation $\sigma=0.01$ m/s. The position measurement noise has standard deviation $\sigma=0.5$ m, and the velocity measurement noise has standard deviation $\sigma=0.2$ m/s. The process and measurement noise covariance matrices $Q$ and $R$ are set accordingly. The initial state covariance matrix is set as $P_{-N}^{-1}=0$, which is a common selection [46, 47]. The set of possible states at simulation time step $k=0$ is $\mathcal{X}_{k=0}=\mathcal{X}_{p0}\times\mathcal{X}_{v0}$ where $\mathcal{X}_{p0}$ and $\mathcal{X}_{v0}$ are again regular hexagons. The initial position state set has inner radius 2 m and is centered at $(x,y)=(-4,1)$. The initial velocity state set has inner radius 1 m/s and is centered at $(\dot{x},\dot{y})=(0,0)$. The true initial state is $\mathbf{x}_{0}=\begin{bmatrix}-5&2&-0.6&0.2\end{bmatrix}^{T}$. The moving horizon length is $N=15$. For simulation time steps less than $k=N$, the horizon length is set to $k$, and $\mathcal{X}_{-N}=\mathcal{X}_{k=0}$. A pre-determined sequence of inputs $\mathbf{u}_{k}$ is applied to the system. Simulated noise inputs are sampled from the previously described truncated normal distributions. The simulation length is 40 time-steps.
@@
Simulated states, MHE estimates, and feasible sets for the numerical example are given in Fig. 5, with Algorithm 5 used to compute the MHE estimates and feasible sets. The left sub-figure depicts the position results while the right sub-figure depicts the velocity results. The RMS position error from the measurements is 0.505 m and the RMS velocity error is 0.200 m/s. The RMS position error from the MHE solution is 0.224 m and the RMS velocity error is 0.066 m/s. Using the computer described in Fig. 4, the average and maximum solution times for this numerical example are ${2.6}\mathrm{e}{-4}$ s and ${5.7}\mathrm{e}{-4}$ s, respectively. The true state and MHE estimate always lie within the feasible set while the measurement may lie outside the feasible set. In this example, the widths of the feasible sets are generally much larger than the MHE RMS errors. By combining SVSE and MHE as has been done here, one can compute an optimal state estimate in the sense of (30) while simultaneously computing the feasible set $\mathcal{X}_{0}$ that contains all possible states of the system.
@@ VI-D Safety Verification
In this example, the discrete time step is $\Delta t=0.5$ s. The state feedback gain $K$ is computed as the discrete LQR gain with quadratic cost function matrices $Q=\text{diag}(\begin{bmatrix}1&1&0&0\end{bmatrix})$ and $R=0.1I$. The set of feasible states $\mathcal{S}$ is the same as that presented in Sec. VI-C. The disturbance set $\mathcal{W}$ is defined as $\mathcal{W}=\mathcal{W}_{p}\times\mathcal{W}_{v}$ with $\mathcal{W}_{p}$ and $\mathcal{W}_{v}$ regular hexagons of radius 0.01 m and 0.2 m/s, respectively. The center of $\mathcal{W}_{v}$ is $(\dot{x},\dot{y})=(0,0.5)$ while $\mathcal{W}_{p}$ is centered at $\mathbf{0}$. The initial state set $\mathcal{X}_{0}$ is given as $\mathcal{X}_{0}=\mathcal{X}_{p0}\times\mathcal{X}_{v0}$ where $\mathcal{X}_{p0}$ and $\mathcal{X}_{v0}$ are regular hexagons of radius 0.5 m and 0.5 m/s, respectively, and with centers $(x,y)=(1,0)$ and $(\dot{x},\dot{y})=(0,0)$. The infeasibility checking interval in Algorithm 1 is set to $k_{\text{inf}}=1$.
The results for this numerical example are shown in Fig. 6. Algorithm 1 finds a certificate of infeasibility for $\mathcal{X}_{k}\cap_{\begin{bmatrix}I&0\end{bmatrix}}\mathcal{O}$ for all time steps $k$, indicating that trajectories of the system are guaranteed to be safe. The sets $\mathcal{X}_{k}$ are color coded to show how many ADMM iterations were required to generate the infeasibility certificate. In the majority of cases, only a single iteration is required. Using the desktop computer described in Fig. 4, the average and maximum times to produce the infeasibility certificates are ${6.1}\mathrm{e}{-5}$ s and ${1.3}\mathrm{e}{-4}$ s, respectively.
@@
VII Conclusion
@@
This paper considers the problem of solving QPs for which the constraints are represented as constrained zonotopes. Novel reachability computations were proposed that produce constrained zonotopes for which the generator and constraint matrices have much higher sparsity than existing methods. A structure-exploiting ADMM algorithm was proposed for solving the QPs. The numerical results demonstrate that when using the sparse reachability calculations to build an MPC problem, the ADMM algorithm solves the resulting QP faster than state-of-the-art QP solvers and a sparse MPC problem formulation, with the solution time gap increasing as the MPC horizon increases. Additionally, reachability computations were used to build a state estimator that combines MHE with SVSE. Efficient safety verification of system trajectories using ADMM infeasibility detection was demonstrated as well.
@@
This paper highlights that reachability calculations using zonotopes and constrained zonotopes can be used to build optimal control and estimation problems with exploitable structure. To facilitate this workflow, the open-source toolbox ZonoOpt was developed.

### v2

@@ VI-C Moving Horizon Estimation and Set-Valued State Estimation
Consider the combined MHE and SVSE problem described in Sec. V-B with a noise-perturbed double integrator model where $A$, $B$, $\mathbf{x}_{k}$, and $\mathbf{u}_{k}$ are defined as in (35) and have the same dimensions as in Sec. VI-B. The measurement matrix is $C=I$. The discrete time step is $\Delta t=1$ s. The set of possible states is given as $\mathcal{S}=\mathcal{S}_{p}\times\mathcal{S}_{v}$ where $\mathcal{S}_{p}$ and $\mathcal{S}_{v}$ are both regular hexagons. The inner radii of $\mathcal{S}_{p}$ and $\mathcal{S}_{v}$ are 500 m and 1 m/s, respectively. The process and measurement noise similarly have the forms $\mathcal{W}=\mathcal{W}_{p}\times\mathcal{W}_{v}$ and $\mathcal{V}=\mathcal{V}_{p}\times\mathcal{V}_{v}$ where $\mathcal{W}_{p}$, $\mathcal{V}_{p}$, $\mathcal{W}_{v}$, and $\mathcal{V}_{v}$ are regular hexagon approximations of circular bounds on position and velocity noise distributions. Truncated normal distributions are used such that the position and velocity noise sets have inner radii of $2\sigma$. Zero-mean noise is used in all cases. The position process noise has standard deviation $\sigma=0.001$ m, and the velocity process noise has standard deviation $\sigma=0.01$ m/s. The position measurement noise has standard deviation $\sigma=0.5$ m, and the velocity measurement noise has standard deviation $\sigma=0.2$ m/s. The process and measurement noise covariance matrices $Q$ and $R$ are set accordingly. The initial state covariance matrix is set as $P_{-N}^{-1}=0$, which is a common selection [53, 54]. The set of possible states at simulation time step $k=0$ is $\mathcal{X}_{k=0}=\mathcal{X}_{p0}\times\mathcal{X}_{v0}$ where $\mathcal{X}_{p0}$ and $\mathcal{X}_{v0}$ are again regular hexagons. The initial position state set has inner radius 2 m and is centered at (x,y)=(4,1)(x,y)=(-4,1). The initial velocity state set has inner radius 1 m/s and is centered at $(\dot{x},\dot{y})=(0,0)$. The true initial state is $\mathbf{x}_{0}=\begin{bmatrix}-5&2&-0.6&0.2\end{bmatrix}^{T}$. The moving horizon length is $N=15$. For simulation time steps less than $k=N$, the horizon length is set to $k$, and $\mathcal{X}_{-N}=\mathcal{X}_{k=0}$. A pre-determined sequence of inputs $\mathbf{u}_{k}$ is applied to the system. Simulated noise inputs are sampled from the previously described truncated normal distributions. The simulation length is 40 time-steps.
@@
Simulated states, MHE estimates, and feasible sets for the numerical example are given in Fig. 6, with Algorithm 5 used to compute the MHE estimates and feasible sets. The left sub-figure depicts the position results while the right sub-figure depicts the velocity results. The RMS position error from the measurements is 0.505 m and the RMS velocity error is 0.200 m/s. The RMS position error from the MHE solution is 0.224 m and the RMS velocity error is 0.066 m/s. Using the computer described in Fig. 5, the average and maximum solution times for this numerical example are $0.026$ ms and $0.057$ ms, respectively. The true state and MHE estimate always lie within the feasible set while the measurement may lie outside the feasible set. In this example, the widths of the feasible sets are generally much larger than the MHE RMS errors. By combining SVSE and MHE as has been done here, one can compute an optimal state estimate in the sense of (31) while simultaneously computing the feasible set $\mathcal{X}_{0}$ that contains all possible states of the system.
@@ VI-D Safety Verification
[…]
The results for this numerical example are shown in Fig. 7. Algorithm 1 finds a certificate of infeasibility for $\mathcal{X}_{k}\cap_{\begin{bmatrix}I&0\end{bmatrix}}\mathcal{O}$ for all time steps $k$, indicating that trajectories of the system are guaranteed to be safe. The sets $\mathcal{X}_{k}$ are color coded to show how many ADMM iterations were required to generate the infeasibility certificate. In the majority of cases, only a single iteration is required. Using the desktop computer described in Fig. 5, the average and maximum times to produce the infeasibility certificates are $0.0061$ ms and $0.013$ ms, respectively.
@@
VII Experimental Results
[…]
Figure 8: Experimental implementation of combined MHE and SVSE for robot localization. Feasible sets $\mathcal{X}_{k}$, motion capture positions $\mathbf{x}_{p,k}$, measurements $\mathbf{y}_{p,k}$, and MHE estimates $\hat{\mathbf{x}}_{p,k}$ are overlaid on top of corrected images from an overhead fisheye camera. Due to imperfect correction of the fisheye effect and differences in sampling rate, images of the robot do not perfectly coincide with the motion capture positions $\mathbf{x}_{p,k}$.
[…]
The ROSbot 2R is a differential drive robot described by the equations of motion
$v=\frac{d}{4}(\omega_{R}+\omega_{L})\;,\;\omega=\frac{d}{2l}(\omega_{R}-\omega_{L})\;,$ (37a)
$\dot{x}=v\cos{(\theta)}\;,\;\dot{y}=v\sin{(\theta)}\;,\;\dot{\theta}=\omega\;,$ (37b)
where $v$ is linear velocity, $\omega$ is angular velocity, $\theta$ is the heading angle, and $x$ and $y$ are position coordinates. The angular velocities of the right and left wheels are $\omega_{R}$ and $\omega_{L}$, respectively, $d$ is the wheel diameter, and $l$ is the distance between the right and left wheels. Note that (37b) is the unicycle model, which is differentially flat in $x$ and $y$ and can be described using the double integrator model [46]. As such, the SVSE/MHE uses the discrete-time double integrator model (35).
The control inputs for this model are the lab frame acceleration components $\mathbf{u}=\begin{bmatrix}\tilde{a}_{x}&\tilde{a}_{y}\end{bmatrix}^{T}$. For the experiment, we define the measurements to be the lab frame position and velocity components $\mathbf{y}=\begin{bmatrix}\tilde{x}&\tilde{y}&\tilde{v}_{x}&\tilde{v}_{y}\end{bmatrix}^{T}$. The measurement matrix is $C=I$. The tilde notation is used here to indicate a measured quantity. The inputs and measurements are derived from motion capture position feedback, wheel motor encoder feedback, and IMU feedback.
The IMU directly senses body frame acceleration $\tilde{\mathbf{a}}^{b}=\begin{bmatrix}\tilde{a}^{b}_{x}&\tilde{a}^{b}_{y}\end{bmatrix}^{T}$ and angular velocity. The IMU also outputs an orientation quaternion based on the integrated angular velocity. The robot’s heading angle in the lab frame, $\tilde{\theta}$, is derived from this orientation quaternion.
The wheel motor encoders provide wheel angular velocity feedback. The two right wheel angular velocities are averaged to give $\tilde{\omega}_{R}$, and the two left wheel angular velocities are averaged to give $\tilde{\omega}_{L}$. The body frame velocity measurement is then computed as $\tilde{\mathbf{v}}^{b}=\begin{bmatrix}\tilde{v}&0\end{bmatrix}^{T}$ where $\tilde{v}$ is computed from $\tilde{\omega}_{R}$ and $\tilde{\omega}_{L}$ using (37a).
The motion capture system provides high-accuracy position feedback signals $\mathbf{x}_{p}=\begin{bmatrix}x&y\end{bmatrix}^{T}$. To simulate less precise position feedback such as that which might be provided by GPS, random noise drawn from a $\pm 2\sigma$-bounded Gaussian distribution was added to the motion capture feedback. These corrupted position feedback signals are denoted as $\tilde{\mathbf{x}}_{p}=\begin{bmatrix}\tilde{x}&\tilde{y}\end{bmatrix}^{T}$, and the position measurement uncertainty standard deviation is given as $\sigma_{p}$.
Transforming body frame quantities into the lab frame, the input and measurement vectors used in the SVSE/MHE are
$\mathbf{u}=\begin{bmatrix}\tilde{a}_{x}\\ \tilde{a}_{y}\end{bmatrix}=\begin{bmatrix}\cos(\tilde{\theta})&-\sin(\tilde{\theta})\\ \sin(\tilde{\theta})&\cos(\tilde{\theta})\end{bmatrix}\begin{bmatrix}\tilde{a}^{b}_{x}\\ \tilde{a}^{b}_{y}\end{bmatrix}\;$, (38a)
$\mathbf{y}=\begin{bmatrix}\tilde{x}\\ \tilde{y}\\ \tilde{v}_{x}\\ \tilde{v}_{y}\end{bmatrix}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&\cos(\tilde{\theta})&-\sin(\tilde{\theta})\\ 0&0&\sin(\tilde{\theta})&\cos(\tilde{\theta})\end{bmatrix}\begin{bmatrix}\tilde{x}\\ \tilde{y}\\ \tilde{v}\\ 0\end{bmatrix}\;$. (38b)
[…]
The acceleration uncertainty set in the lab frame is given as
$\mathcal{W}_{u}=\begin{bmatrix}\cos(\tilde{\theta})&-\sin(\tilde{\theta})\\ \sin(\tilde{\theta})&\hphantom{-}\cos(\tilde{\theta})\end{bmatrix}\mathcal{W}_{u}^{b}\oplus\mathcal{E}_{u}\;,$ (39)
where $\mathcal{W}_{u}^{b}$ is the acceleration measurement error set in the body frame and $\mathcal{E}_{u}$ accounts for heading angle uncertainty. In this case, $\mathcal{W}_{u}^{b}\equiv[\mathbf{e}_{u}^{b}]$ where $\mathbf{e}_{u}^{b}]=\begin{bmatrix}[e_{a,x}^{b}]&[e_{a,y}^{b}]\end{bmatrix}^{T}$ is a box, and $e^{b}_{a,x}]$ and $e^{b}_{a,y}]$ are the body frame acceleration error intervals. An over-approximation of the uncertainty in $\mathcal{W}_{u}$ due to heading angle uncertainty is
$\mathcal{E}_{u}=\begin{bmatrix}[e_{c\theta}][e_{a,x}^{b}]-[e_{s\theta}][e_{a,y}^{b}]\\ [e_{s\theta}][e_{a,x}^{b}]+[e_{c\theta}][e_{a,y}^{b}]\end{bmatrix}\;$. (40)
To implement (39), $\mathcal{W}_{u}^{b}$ and $\mathcal{E}_{u}$ were first converted to zonotopes, and then the identities in (3) were used to compute $\mathcal{W}_{u}$. The disturbance set $\mathcal{W}$ is then given as $\mathcal{W}=B\mathcal{W}_{u}$.
The measurement error set is given as $\mathcal{V}=\mathcal{V}_{p}\times\mathcal{V}_{v}$, where $\mathcal{V}_{p}$ is the position measurement error set and $\mathcal{V}_{v}$ is the velocity measurement error set. The position measurement error set is
$\mathcal{V}_{p}=\begin{bmatrix}\sigma_{p}&0\\ 0&\sigma_{p}\end{bmatrix}\mathcal{Z}_{\circ}\;,$ (41)
where $\mathcal{Z}_{\circ}$ is an outer approximation of a unit circle, taken here as an eight-sided zonotope. The velocity measurement error set $\mathcal{V}_{v}$ was computed analogously to $\mathcal{W}_{u}$, with the body frame velocity error set given as $\mathcal{V}^{b}_{v}\equiv[\mathbf{e}_{v}^{b}]$ where $\mathbf{e}_{v}^{b}]=\begin{bmatrix}[-e_{v},e_{v}]&[-\epsilon,\epsilon]\end{bmatrix}^{T}$. In this case, $e_{v}$ is the maximum error in the velocity derived from the wheel motor encoder measurements, and $0<\epsilon\ll 1$.
For the MHE, covariance matrices $Q$ and $R$ were estimated by transforming body frame covariance matrices into the lab frame as required for velocity and acceleration. Standard deviations are taken to be one half of the uncertainty bound, e.g., the body frame forward velocity standard deviation is $\sigma_{v}=\frac{1}{2}e_{v}$. The covariance matrices do not account for heading angle uncertainty contributions. As in Sec. VI-C, the initial inverse covariance matrix was taken to be $P^{-1}_{N}=0$.
The set of possible states is $\mathcal{S}=\mathcal{S}_{p}\times\mathcal{S}_{v}$, where $\mathcal{S}_{p}$ and $\mathcal{S}_{v}$ are the possible position and velocity states. The possible position states are defined by the dimensions of the laboratory space, and the possible velocity states are given as
$\mathcal{S}_{v}=\begin{bmatrix}v_{\text{max}}&0\\ 0&v_{\text{max}}\end{bmatrix}\mathcal{Z}_{\circ}\;,$ (42)
where $v_{\text{max}}$ is a bound on the translational velocity.
Position feedback was published at a rate of $100$ Hz, and the robot published both the IMU and wheel motor encoder feedback messages at $20$ Hz. The SVSE/MHE ran at a rate of $10$ Hz and was triggered off of the IMU message, i.e., the SVSE/MHE algorithm executed immediately following receipt of the IMU message, every other IMU message. The horizon length was $N=20$. The initial feasible set was computed from the initial measurements as
$\mathcal{X}_{0}=\begin{bmatrix}\tilde{x}_{0}&\tilde{y}_{0}&\tilde{v}_{x,0}&\tilde{v}_{y,0}\end{bmatrix}^{T}\oplus\mathcal{V}\;.$ (43)
To mitigate complexity growth of the initial feasible set $\mathcal{X}_{-N}$ over repeated SVSE/MHE iterations, $\mathcal{X}_{-N}$ was periodically replaced with its bounding box approximation. The bounding box was computed by calculating the support, $\text{sup}(\mathcal{X}_{-N},\mathbf{d})=\max_{\mathbf{z}\in\mathcal{X}_{-N}}\langle\mathbf{z},\mathbf{d}\rangle$, along each axis-aligned direction $\mathbf{d}$. The resulting box is then converted to a zonotope. Here, each bounding box calculation requires eight invocations of Algorithm 1. The factorizations of $M$ and $AA^{T}$ were reused between optimizations for efficiency. Bounding box approximations were computed every 10 SVSE/MHE iterations following execution of Algorithm 3.
TABLE II: Parameters for SVSE/MHE experimental implementation.
Param. $d$ $v_{\text{max}}$ $\sigma_{p}$ $e_{\theta}$ $e_{v}$ $e^{b}_{u,x}$ $e^{b}_{u,y}$
Units m m/s m rad m/s m/s2 m/s2
Value 0.084 1.1 0.3 0.262 0.25 3.0 3.0
Parameters for the experimental implementation are given in Table II. The heading uncertainty $e_{\theta}$ accounts for both initialization error and errors in the IMU’s heading computation. The velocity uncertainty $e_{v}$ is dominated by errors due to asynchronous sampling between the IMU and wheel motor encoder measurements. Acceleration uncertainty bounds were estimated from inspection of recorded acceleration data.
Results for this experiment are displayed in Fig. 8. Here, the MHE position estimates $\hat{\mathbf{x}}_{p,k}$ and projections of the SVSE bounding sets $\mathcal{X}_{k}$ onto the position states are overlaid on top of overhead camera images of the laboratory space. The motion capture positions are given by $\mathbf{x}_{p,k}$ and are taken to be ground truth. The position measurement feedback for the SVSE/MHE is given by $\mathbf{y}_{p,k}$. Because there is significant overlap between adjacent $\mathcal{X}_{k}$, only every tenth iteration is plotted. Effectiveness of the SVSE is verified by the observation that $\mathbf{x}_{p,k}\in[I\;0]\mathcal{X}_{k}\;\forall k$. The MHE significantly improved position estimation accuracy when compared to the position measurements. The RMS error for the position measurements $\mathbf{y}_{p,k}$ was 0.31 m, while the RMS error for the MHE position estimates $\hat{\mathbf{x}}_{p,k}$ was 0.11 m.
VIII Conclusion
@@
This paper considers the problem of solving QPs for which the constraints are represented as constrained zonotopes. Novel reachability computations were proposed that produce constrained zonotopes for which the generator and constraint matrices have much higher sparsity than existing methods. An ADMM algorithm was proposed to efficiently solve the QPs. The numerical results demonstrate that when using the sparse reachability calculations to build an MPC problem, the ADMM algorithm solves the resulting QP faster than state-of-the-art QP solvers and a sparse MPC problem formulation, with the solution time gap increasing as the MPC horizon increases. Additionally, reachability computations were used to build a state estimator that combines MHE with SVSE, and this state estimator was experimentally applied for online robot localization. Efficient safety verification of system trajectories using ADMM infeasibility detection was demonstrated as well.
@@
This paper highlights that reachability calculations with zonotopes and constrained zonotopes can be used to build optimal control and estimation problems amenable to efficient optimization. To facilitate workflows based on real-time reachability analysis and optimization, the open-source toolbox ZonoOpt was developed.